OpenStudy (anonymous):

need help solving with elimination method. y=4x-1 y=2/3x-1

OpenStudy (aravindg):

subtract first and second equation!!

OpenStudy (aravindg):

@channie2003 do it and you will get x !

OpenStudy (anonymous):

thank you im just starting a new class and trying to get back into how to do these things I am extremely horrible at math. I get confuse very easy

OpenStudy (hba):

Practice makes perfect :)

OpenStudy (aravindg):

every warrior was once a crying baby :P

OpenStudy (texaschic101):

y = 4x - 1 y = 2/3x - 1 ------------ Lets get rid of the fraction in the second equation by multiplying it by 3 (3)y = (3)2/3x - (3)1 3y = 2x - 3 Now your problem reads.... y = 4x - 1 3y = 2x - 3---> (-2)3y = (-2)2x - (-2)3 -------------- y = 4x - 1 -6x = -4x + 6 (result of multiplying by -2) -------------add -5y = 5 (as you can see, the 4x's cancel each other out) -5y/-5 = 5/-5 (divide by -5 to get x by itself) y = -1 now sub -1 into either of the first two equations to find x. y = 4x - 1 -1 = 4x - 1 -1 + 1 = 4x -1 + 1 (add 1 to both sides to get 4x by itself) 0 = 4x 0/4 = x 0 = x check.... y = 4x - 1 -1 = 4(0) -1 -1 = -1 (correct) y = 2/3x - 1 -1 = 2/3(0) - 1 -1 = -1 (correct) ANSWER : x = 0, y = -1